### v1

@@ Abstract
Various natural and engineered systems, from urban traffic flow to the human brain, have been described by large-scale networked dynamical systems. Despite their vast differences, these systems are often similar in being comprised of numerous microscopic subsystems with complex nonlinear dynamics and interactions that give rise to diverse emergent macroscopic behaviors. As such, a long-standing question across various fields has been to understand why and how various forms of macroscopic behavior emerge from underlying microscopic dynamics. Motivated by a growing body of empirical observations, in this work we focus on linearity as one of the most fundamental aspects of system dynamics, and develop a general theoretical framework for the interplay between spatial averaging, decaying microscopic correlations, and emergent macroscopic linearity. Using and extending the theory of mixing sequences, we show that in a broad class of autonomous nonlinear networked systems, the dynamics of the average of all subsystems’ states becomes asymptotically linear as the number of subsystems grows to infinity, provided that (in addition to technical assumptions) pairwise correlations between subsystems decay to 0 as their pairwise distance grows to infinity. We prove this result when the latter distance is between subsystems’ linear indices or spatial locations, and provide extensions to linear time-invariant (LTI) limit dynamics, finite-sample analysis of rates of convergence, and networks of spatially-embedded subsystems with random locations. To our knowledge, this work is the first rigorous analysis of macroscopic linearity in large-scale heterogeneous networked systems, and provides a solid foundation for further theoretical and empirical analyses in various domains of science and engineering.
I Introduction
@@ II Notation
III Preliminaries: Mixing Sequences
@@
We are now ready to present the following result, which is a generalization of the standard CLT for i.i.d. sequences [33, Thm 3.4.1] to $\rho$-mixing sequences of random variables.
Proposition III.3.
(CLT for $\rho$-mixing sequences [35, Thm B]). Consider a $\rho$-mixing sequence of random variables $\xi_{1},\xi_{2},$.... and define its cumulative sum and cumulative variance as
@@
It is instructive to compare Proposition III.3 with the standard CLT for i.i.d. sequences [33, Thm 3.4.1]. Both versions assume that all $\xi_{i}$ have finite mean and variance; they slightly differ in that Proposition III.3 assumes $\xi_{i}$’s are already mean-subtracted while the standard version mean-subtracts the sum. Condition (4b) asks that the ratio of the variance of shifted partial sums to that of initial partial sums remains bounded. This is a generalization of the ‘identically distributed’ assumption in the standard version, in which case the left hand side of (4b) is always 1. Finally, $\rho$-mixingness is a generalization of independence, as noted earlier. Importantly, however, condition (4c) ensures that the sequence $\{\xi_{i}\}$ is not any $\rho$-mixing sequence, but one in which the correlations between neighboring elements decays sufficiently fast. This is a subtle point and we will get back to it later in Section IV when we define residual factors. For now, note the normalization by $\sigma_{N}$ in (5) vs. the standard normalization by $\sqrt{N}$ for i.i.d. sequences. This difference stems from the fact that the growth rate of the cumulative variance ($\sigma_{N}^{2}$) of $\rho$-mixing sequences can vary, whereas $\sigma_{N}^{2}$ always grows as $N$ for i.i.d. sequences.
In what follows, we will extend Proposition III.3 in a number of ways, and use the results to prove the Gaussianity of joint state-noise distributions under spatial averaging, using which we will prove the linearity of spatially-averaged dynamics.
IV Linearizing Effect of Spatial Averaging on Sequences of Dynamical Systems
IV-A Problem Formulation
@@
In the rest of this section we will first lay the statistical foundation of our framework in Section IV-B, where we extend and further characterize the properties of $\rho$-mixing sequences. In Section IV-C we then use this foundation to present our main result of this section, i.e., the solution to Problem 1. Under additional stationarity assumptions, we further strengthen this result to prove linear time-invariant (LTI) limit dynamics, and finally characterize the rate of this convergence to linearity in Section IV-E for the special case of i.i.d. sequences.
IV-B Multivariate $\rho$-Mixing Sequences
In this subsection we extend Definition III.1 to vector-valued stochastic processes, and then prove some of their properties, including a more practical CLT. These results will play a central role in our proof of the linearizing effect of spatial averaging in Theorem IV.8.
The following is a natural generalization of Definition III.1.
Definition IV.1.
@@ IV-B Multivariate $\rho$-Mixing Sequences
To put this definition in perspective, if $\{\boldsymbol{\xi}_{i}\}_{i=1}^{\infty}$ is i.i.d. then ${\rm Cov}(S_{N})=N{\rm Cov}(\boldsymbol{\xi}_{1})$ and so $h(N)=1$ is a residual factor. Note, also, that residual factors are not unique. In the i.i.d. case, $h(N)=c$ is also a residual factor for any constant $c$. Further, note that Definition IV.2 does not require the limit to be nonzero. Thus, $h(N)=N$, $\log(N)$, and $e^{N}$ are all valid residual factors for an i.i.d. sequence. Finally, and related to the last point, a residual factor always exists for any $\rho$-mixing sequence. The following Theorem formalizes this existence.
Theorem IV.3.
@@
for all $\ell,m=1,\dots,q$. Put together, $\lim_{N\to\infty}\frac{1}{N^{2}}{\rm Cov}(S_{N})=0$, completing the proof. ∎
A corollary to Theorem IV.3 is that any $h(N)\geq N$ is also a valid residual factor for any $\rho$-mixing sequence with uniformly-bounded variance. However, residual factors are practically useful only when they grow at the same rate as ${\rm Cov}(S_{N})$ itself and the limit in (11) is nonzero. This “smallest” residual factor is what we will use in the sequel, even though the forthcoming results will still be correct, but not necessarily useful, for all valid residual factors.
Going back to Proposition III.3, note also the similarity between (4c) and (11), where $h(N)$ in (4c) is clearly a residual factor for $\{\xi_{k}\}_{i=1}^{\infty}$ in Proposition III.3. As we will see later, whether a residual factor of a $\rho$-mixing sequence is slowly-varying remains to be a critical condition for validity of CLT (and in turn asymptotic linearity) for that sequence.
@@
    In other words, the correlation decay rate of $\frac{1}{|i-j|}$ is “sufficiently fast” so that ${\rm Var}(S_{N})$ grows sufficiently slowly and $h(N)$ becomes slowly varying. As we will see in Theorem IV.6, this allows for CLT to hold for this sequence.
@@
    Therefore in this case, unlike case (i), correlations decay too slowly. As a result, ${\rm Var}(S_{N})$ grow too fast and $h(N)$ is not slowly-varying. As we will see in Theorem IV.6, this prevents CLT to hold for this sequence, despite being $\rho$-mixing.
@@
In summary, it follows from Theorem IV.4 that $\rho(n)\propto\frac{1}{n}$ is the rough “boundary” between decay rates that are sufficiently fast and those that are too slow.
@@
where $\rho_{\zeta}(n)$ and $\rho_{\xi}(n)$ are defined as in (1) for the respective sequences $\{\boldsymbol{\zeta}_{i}\}$ and $\{\boldsymbol{\xi}_{i}\}$. The theorem then follows by letting $n\to\infty$ in (20) and using the mixing property of $\{\boldsymbol{\xi}_{i}\}$. ∎
The next result extends and improves the CLT in Proposition III.3 to the case of multivariate $\rho$-mixing sequences.
Theorem IV.6.
@@
Proof.
Let $\boldsymbol{\theta}\neq\mathbf{0}$ be an arbitrary constant vector. By Lemma IV.5, the sequence of random variables $\{\boldsymbol{\theta}^{T}\boldsymbol{\xi}_{i}\}_{i\geq 1}$ is $\rho$-mixing. Define the cumulative sum and cumulative variance of this sequence as
@@
The sequence $\{\boldsymbol{\theta}^{T}\boldsymbol{\xi}_{i}\}_{i\geq 1}$ follows (4a) by assumption. It also follows (4b) due to (21). We need to show that it also satisfies (4c) before we can use the CLT in Proposition III.3.
@@
is slowly varying, and thus $\sigma_{N}^{2}$ satisfies (4c). Therefore, Proposition III.3 holds and, from (5) and (26), as $N\to\infty$
@@
Therefore, (29) holds for all $\boldsymbol{\theta}\in^{q}$. According to Cramer-Wold device theorem [33, Thm 3.9.5], it then follows that
@@
completing the proof. ∎
Theorem IV.6 provides the foundation for the asymptotic analysis of spatially averaged dynamics formulated in Problem 1, which we undertake next.
IV-C Asymptotic Linearity Under Spatial Averaging
In this section we use the results in Section IV-B to tackle Problem 1. Throughout this section, we make the following assumptions on the population dynamics (6). For ease of notation, for all $i,t$ let
@@ IV-C Asymptotic Linearity Under Spatial Averaging
The key assumption in Assumption IV.7 is the $\rho$-mixing assumption in (A5) which, in essence, prevents global synchrony. In other words, it prevents a case where all $\{\mathbf{y}_{i}(t)\}_{i=1}^{\infty}$ are strongly correlated with each other, in which case averaging does not do much—the average of $\{\mathbf{y}_{i}(t)\}_{i=1}^{\infty}$ becomes very similar to any one of them.
The remaining 4 assumptions are technical and automatically satisfied in most practical applications. Assumptions (A1) and (A2) are parallel to (4a) and (4b) in Proposition III.3 and described thereafter. Next, assumption (A3) asks for lack of instantaneous effect, where each noise variable takes at least one time step before affecting the state. Finally, note that Assumption (A4) is practically equivalent to uniform boundedness of state-noise trajectories. If $(\mathbf{x}_{\mathcal{N}_{i}},\mathbf{w}_{i})$ are guaranteed to remain within a compact region $\mathcal{D}$, as is the case for any real-world state/noise variable, then $f_{i}$ needs to be bounded only on $\mathcal{D}$ and its value can be replaced by an arbitrary finite value (say, 0) outside of $\mathcal{D}$ without affecting the solutions of (6). This is true, e.g., if $f_{i}$ is any analytic function on $\mathcal{D}$.
@@
Proof.
Fix $t$, and let $\mathbf{z}_{i}(t)$ be as in (32). From (6), Assumption (A5), and Lemma IV.5, the sequence $\{\mathbf{z}_{i}(t)\}_{i=1}^{\infty}$ is also $\rho$-mixing. Since mean-centering preserves correlation coefficients, $\{\mathbf{z}_{i}(t)-\mathbb{E}[\mathbf{z}_{i}(t)]\}_{i=1}^{\infty}$ is $\rho$-mixing as well. The latter sequence is mean-zero by construction and finite-variance by Assumptions (A4) and (A1). Furthermore, $h(N)$ is a residual factor for both $\{\mathbf{x}_{i}(t+1)\}_{i=1}^{\infty}$ and $\big\{\begin{bmatrix}\mathbf{x}_{i}(t)^{T}&\mathbf{w}_{i}(t)^{T}\end{bmatrix}^{T}\big\}_{i=1}^{\infty}$, so both of the limits
@@
exist. Since the limit of the covariance between $\bar{\mathbf{x}}(t+1)$ and $\bar{\mathbf{y}}(t)$ also exists by assumption, $h(N)$ is a (slowly-varying) residual factor for $\{\mathbf{z}_{i}(t)-\mathbb{E}[\mathbf{z}_{i}(t)]\}_{i=1}^{\infty}$. These, together with Assumption (A2), satisfy all the requirements of Theorem IV.6. Therefore,
@@
Since $\mathbf{z}_{\infty}$ is normally distributed, the right hand side in (IV-C) equals
@@
which simplifies to $\mathbf{A}_{\infty}(t)\boldsymbol{\xi}+\mathbf{B}_{\infty}(t)\boldsymbol{\omega}$ in (IV.8) because of the independence of $\bar{\mathbf{x}}_{\infty}(t)$ and $\bar{\mathbf{w}}_{\infty}(t)$ resulting from Assumption (A3). ∎
As far as linearity of the limit dynamics is concerned, it is important to note that the matrices $\mathbf{A}_{\infty}(t)$ and $\mathbf{B}_{\infty}(t)$ in Theorem IV.8 depend on the distributions of states $\{\mathbf{x}_{i}(t)\}_{i=1}^{\infty},\{\mathbf{x}_{i}(t+1)\}_{i=1}^{\infty}$ and noises $\{\mathbf{w}_{i}(t)\}_{i=1}^{\infty}$, but not on their specific values (realizations). In other words, for any choice of model in (6), i.e., for any choice of $\{f_{i}(\cdot)\}_{i=1}^{\infty}$, $p_{0}$, and $p_{w}$, the matrices $\mathbf{A}_{\infty}(t)$ and $\mathbf{B}_{\infty}(t)$ can be in theory pre-computed for all $t\geq 0$, and would be the same for all realizations of noises and states generated by this system. This is important for the dynamics to be actually linear, as illustrated by the following simple example.
Example IV.9.
@@
IV-D Time-Invariant Limit Dynamics

### v2

@@ Abstract
Various natural and engineered systems, from urban traffic flow to the human brain, can be described by large-scale networked dynamical systems. These systems are similar in being comprised of a large number of microscopic subsystems, each with complex nonlinear dynamics and interactions, that collectively give rise to different forms of macroscopic dynamics. Despite significant research, why and how various forms of macroscopic dynamics emerge from underlying micro-dynamics remains largely unknown. In this work we focus on linearity as one of the most fundamental aspects of system dynamics. By extending the theory of mixing sequences, we show that in a broad class of autonomous nonlinear networked systems, the dynamics of the average of all subsystems’ states becomes asymptotically linear as the number of subsystems grows to infinity, provided that, in addition to technical assumptions, pairwise correlations between subsystems decay to 0 as their pairwise distance grows to infinity. We prove this result when the latter distance is between subsystems’ linear indices or spatial locations, and provide extensions to linear time-invariant (LTI) limit dynamics, finite-sample analysis of rates of convergence, and networks of spatially-embedded subsystems with random locations. To our knowledge, this work is the first rigorous analysis of macroscopic linearity in large-scale heterogeneous networked dynamical systems, and provides a solid foundation for further theoretical and empirical analyses in various domains of science and engineering.
keywords:
Large-Scale Complex Systems; Linear/Nonlinear Models; Infinite-Dimensional Systems; Multi-Scale Systems; Statistical Analysis
thanks: A preliminary version of this work appeared at the 61st IEEE Conference on Decision and Control [1]. Corresponding author Erfan Nozari.
1 Introduction
@@ 2 Notation
3 Preliminaries: Mixing Sequences
@@
We are now ready to present the following result, which is a generalization of the standard CLT for i.i.d. sequences [17, Thm 3.4.1] to $\rho$-mixing sequences of random variables.
Proposition 3.3.
(CLT for $\rho$-mixing sequences [25, Thm B]). Consider a $\rho$-mixing sequence of random variables $\xi_{1},\xi_{2},$.... and define its cumulative sum and cumulative variance as
@@
It is instructive to compare Proposition 3.3 with the standard CLT for i.i.d. sequences [17, Thm 3.4.1]. Both versions assume that all $\xi_{i}$ have finite mean and variance; they slightly differ in that Proposition 3.3 assumes $\xi_{i}$’s are already mean-subtracted while the standard version mean-subtracts the sum. Condition (4b) asks that the ratio of the variance of shifted partial sums to that of initial partial sums remains bounded. This is a generalization of the ‘identically distributed’ assumption in the standard version, in which case the left hand side of (4b) is always 1. Finally, $\rho$-mixingness is a generalization of independence, as noted earlier. Importantly, however, condition (4c) ensures that the sequence $\{\xi_{i}\}$ is not any $\rho$-mixing sequence, but one in which the correlations between neighboring elements decays sufficiently fast. This is a subtle point and we will get back to it later in Section 4 when we define residual factors. For now, note the normalization by $\sigma_{N}$ in (5) vs. the standard normalization by $\sqrt{N}$ for i.i.d. sequences. This difference stems from the fact that the growth rate of the cumulative variance ($\sigma_{N}^{2}$) of $\rho$-mixing sequences can vary, whereas $\sigma_{N}^{2}$ always grows as $N$ for i.i.d. sequences.
In what follows, we will extend Proposition 3.3 in a number of ways, and use the results to prove the Gaussianity of joint state-noise distributions under spatial averaging, using which we will prove the linearity of spatially-averaged dynamics.
4 Linearizing Effect of Spatial Averaging on Sequences of Dynamical Systems
4.1 Problem Formulation
@@
In the rest of this section we will first lay the statistical foundation of our framework in Section 4.2, where we extend and further characterize the properties of $\rho$-mixing sequences. In Section 4.3 we then use this foundation to present our main result of this section, i.e., the solution to Problem 1. Under additional stationarity assumptions, we further strengthen this result to prove linear time-invariant (LTI) limit dynamics, and finally characterize the rate of this convergence to linearity in Section 4.5 for the special case of i.i.d. sequences.
4.2 Multivariate $\rho$-Mixing Sequences
In this subsection we extend Definition 3.1 to vector-valued stochastic processes, and then prove some of their properties, including a more practical CLT. These results will play a central role in our proof of the linearizing effect of spatial averaging in Theorem 4.8.
The following is a natural generalization of Definition 3.1.
Definition 4.1.
@@ 4.2 Multivariate $\rho$-Mixing Sequences
To put this definition in perspective, if $\{\boldsymbol{\xi}_{i}\}_{i=1}^{\infty}$ is i.i.d. then ${\rm Cov}(S_{N})=N{\rm Cov}(\boldsymbol{\xi}_{1})$ and so $h(N)=1$ is a residual factor. Note, also, that residual factors are not unique. In the i.i.d. case, $h(N)=c$ is also a residual factor for any constant $c$. Further, note that Definition 4.2 does not require the limit to be nonzero. Thus, $h(N)=N$, $\log(N)$, and $e^{N}$ are all valid residual factors for an i.i.d. sequence. Finally, and related to the last point, a residual factor always exists for any $\rho$-mixing sequence. The following Theorem formalizes this existence.
Theorem 4.3.
@@
for all $\ell,m=1,\dots,q$. Put together, $\lim_{N\to\infty}\frac{1}{N^{2}}{\rm Cov}(S_{N})=0$, completing the proof.
A corollary to Theorem 4.3 is that any $h(N)\geq N$ is also a valid residual factor for any $\rho$-mixing sequence with uniformly-bounded variance. However, residual factors are practically useful only when they grow at the same rate as ${\rm Cov}(S_{N})$ itself and the limit in (11) is nonzero. This “smallest” residual factor is what we will use in the sequel, even though the forthcoming results will still be correct, but not necessarily useful, for all valid residual factors.
Going back to Proposition 3.3, note also the similarity between (4c) and (11), where $h(N)$ in (4c) is clearly a residual factor for $\{\xi_{k}\}_{i=1}^{\infty}$ in Proposition 3.3. As we will see later, whether a residual factor of a $\rho$-mixing sequence is slowly-varying remains to be a critical condition for validity of CLT (and in turn asymptotic linearity) for that sequence.
@@
    In other words, the correlation decay rate of $\frac{1}{|i-j|}$ is “sufficiently fast” so that ${\rm Var}(S_{N})$ grows sufficiently slowly and $h(N)$ becomes slowly varying. As we will see in Theorem 4.6, this allows for CLT to hold for this sequence.
@@
    Therefore in this case, unlike case (i), correlations decay too slowly. As a result, ${\rm Var}(S_{N})$ grow too fast and $h(N)$ is not slowly-varying. As we will see in Theorem 4.6, this prevents CLT to hold for this sequence, despite being $\rho$-mixing.
@@
In summary, it follows from Theorem 4.4 that $\rho(n)\propto\frac{1}{n}$ is the rough “boundary” between decay rates that are sufficiently fast and those that are too slow.
@@
where $\rho_{\zeta}(n)$ and $\rho_{\xi}(n)$ are defined as in (1) for the respective sequences $\{\boldsymbol{\zeta}_{i}\}$ and $\{\boldsymbol{\xi}_{i}\}$. The theorem then follows by letting $n\to\infty$ in (20) and using the mixing property of $\{\boldsymbol{\xi}_{i}\}$.
The next result extends and improves the CLT in Proposition 3.3 to the case of multivariate $\rho$-mixing sequences.
Theorem 4.6.
@@
{proof}
Let $\boldsymbol{\theta}\neq\mathbf{0}$ be an arbitrary constant vector. By Lemma 4.5, the sequence of random variables $\{\boldsymbol{\theta}^{T}\boldsymbol{\xi}_{i}\}_{i\geq 1}$ is $\rho$-mixing. Define the cumulative sum and cumulative variance of this sequence as
@@
The sequence $\{\boldsymbol{\theta}^{T}\boldsymbol{\xi}_{i}\}_{i\geq 1}$ follows (4a) by assumption. It also follows (4b) due to (21). We need to show that it also satisfies (4c) before we can use the CLT in Proposition 3.3.
@@
is slowly varying, and thus $\sigma_{N}^{2}$ satisfies (4c). Therefore, Proposition 3.3 holds and, from (5) and (26), as $N\to\infty$
@@
Therefore, (29) holds for all $\boldsymbol{\theta}\in^{q}$. According to Cramer-Wold device theorem [17, Thm 3.9.5], it then follows that
@@
completing the proof.
Theorem 4.6 provides the foundation for the asymptotic analysis of spatially averaged dynamics formulated in Problem 1, which we undertake next.
4.3 Asymptotic Linearity Under Spatial Averaging
In this section we use the results in Section 4.2 to tackle Problem 1. Throughout this section, we make the following assumptions on the population dynamics (6). For ease of notation, for all $i,t$ let
@@ 4.3 Asymptotic Linearity Under Spatial Averaging
The key assumption in Assumption 4.7 is the $\rho$-mixing assumption in (A5) which, in essence, prevents global synchrony. In other words, it prevents a case where all $\{\mathbf{y}_{i}(t)\}_{i=1}^{\infty}$ are strongly correlated with each other, in which case averaging does not do much—the average of $\{\mathbf{y}_{i}(t)\}_{i=1}^{\infty}$ becomes very similar to any one of them.
The remaining 4 assumptions are technical and automatically satisfied in most practical applications. Assumptions (A1) and (A2) are parallel to (4a) and (4b) in Proposition 3.3 and described thereafter. Next, assumption (A3) asks for lack of instantaneous effect, where each noise variable takes at least one time step before affecting the state. Finally, note that Assumption (A4) is practically equivalent to uniform boundedness of state-noise trajectories. If $(\mathbf{x}_{\mathcal{N}_{i}},\mathbf{w}_{i})$ are guaranteed to remain within a compact region $\mathcal{D}$, as is the case for any real-world state/noise variable, then $f_{i}$ needs to be bounded only on $\mathcal{D}$ and its value can be replaced by an arbitrary finite value (say, 0) outside of $\mathcal{D}$ without affecting the solutions of (6). This is true, e.g., if $f_{i}$ is any analytic function on $\mathcal{D}$.
@@
{proof}
Fix $t$, and let $\mathbf{z}_{i}(t)$ be as in (32). From (6), Assumption (A5), and Lemma 4.5, the sequence $\{\mathbf{z}_{i}(t)\}_{i=1}^{\infty}$ is also $\rho$-mixing. Since mean-centering preserves correlation coefficients, $\{\mathbf{z}_{i}(t)-\mathbb{E}[\mathbf{z}_{i}(t)]\}_{i=1}^{\infty}$ is $\rho$-mixing as well. The latter sequence is mean-zero by construction and finite-variance by Assumptions (A4) and (A1). Furthermore, $h(N)$ is a residual factor for both $\{\mathbf{x}_{i}(t+1)\}_{i=1}^{\infty}$ and $\big\{\begin{bmatrix}\mathbf{x}_{i}(t)^{T}&\mathbf{w}_{i}(t)^{T}\end{bmatrix}^{T}\big\}_{i=1}^{\infty}$, so both of the limits
@@
exist. Since the limit of the covariance between $\bar{\mathbf{x}}(t+1)$ and $\bar{\mathbf{y}}(t)$ also exists by assumption, $h(N)$ is a (slowly-varying) residual factor for $\{\mathbf{z}_{i}(t)-\mathbb{E}[\mathbf{z}_{i}(t)]\}_{i=1}^{\infty}$. These, together with Assumption (A2), satisfy all the requirements of Theorem 4.6. Therefore,
@@
Since $\mathbf{z}_{\infty}$ is normally distributed, the right hand side in (4.3) equals
@@
which simplifies to $\mathbf{A}_{\infty}(t)\boldsymbol{\xi}+\mathbf{B}_{\infty}(t)\boldsymbol{\omega}$ in (4.8) because of the independence of $\bar{\mathbf{x}}_{\infty}(t)$ and $\bar{\mathbf{w}}_{\infty}(t)$ resulting from Assumption (A3).
As far as linearity of the limit dynamics is concerned, it is important to note that the matrices $\mathbf{A}_{\infty}(t)$ and $\mathbf{B}_{\infty}(t)$ in Theorem 4.8 depend on the distributions of states $\{\mathbf{x}_{i}(t)\}_{i=1}^{\infty},\{\mathbf{x}_{i}(t+1)\}_{i=1}^{\infty}$ and noises $\{\mathbf{w}_{i}(t)\}_{i=1}^{\infty}$, but not on their specific values (realizations). In other words, for any choice of model in (6), i.e., for any choice of $\{f_{i}(\cdot)\}_{i=1}^{\infty}$, $p_{0}$, and $p_{w}$, the matrices $\mathbf{A}_{\infty}(t)$ and $\mathbf{B}_{\infty}(t)$ can be in theory pre-computed for all $t\geq 0$, and would be the same for all realizations of noises and states generated by this system. This is important for the dynamics to be actually linear, as illustrated by the following simple example.
Example 4.9.
@@
4.4 Time-Invariant Limit Dynamics